OpenStudy (anonymous):

Find the number of halve in 6/7?

OpenStudy (anonymous):

why 12/7?

OpenStudy (anonymous):

Half of 6/7? Half is equal to 1/2 6/7 * 1/2 = 6/14 = 3/7

OpenStudy (anonymous):

actually 3/7 sorry

OpenStudy (anonymous):

haha :P

OpenStudy (anonymous):

yes:) check : 6over7 is 0.85714... and 3 over 7 is 0.4285... so it halved! 8)